OpenStudy (anonymous):

how to find the area of each section of the trapezoid with no hight or total area giving

OpenStudy (anonymous):

What information do you know?

OpenStudy (anonymous):

The top base is 10 the bottom base is 21 and the side of the triangle is .... I don't know the name of this symbol

OpenStudy (anonymous):

theta? \[\theta\]

OpenStudy (anonymous):

square root symbol 45 in

OpenStudy (anonymous):

So \[\sqrt{45}\] is the length of the side?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok. Are you familiar with the Pythagorean Theorem? \[a ^{2}+b ^{2}=c ^{2}\] ?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

Ok. If you have a right triangle, if you say that c is the hypotenuse (the longest side) and the other two sides are a and b, if you know any of the two sides you can find the third. So for example if you knew what c and a were you could find b by solving \[b=\sqrt{c ^{2}-a ^{2}}\]

OpenStudy (anonymous):

OK, c is square root 45 and a is 5.5

OpenStudy (anonymous):

so what do I do now

OpenStudy (anonymous):

Well, if a=5.5 (which is true if your trapezoid is symmetric), then you just use the formula I put up a second ago. The square of root(45) is just 45, and 5.5^2 is 30.25. You can do the rest.

OpenStudy (anonymous):

b=sqare root 45 to the second power - 5.5 to the second power

OpenStudy (anonymous):

b = square root 2025-30.25 b=square root 1994.75 b=44.66

OpenStudy (anonymous):

is this correct

OpenStudy (anonymous):

No, the square root and the square operators cancel out. \[\sqrt{45}^{2}=45\]

OpenStudy (anonymous):

so it's b=45-5.5 to the second power

OpenStudy (anonymous):

You mean 45-5.5^2? That is b^2. You have to take the square root of that in order to get b.

OpenStudy (anonymous):

yes that is what I meant 5.5^2

OpenStudy (anonymous):

Is it b=45-30.25 b=14.75

OpenStudy (anonymous):

as you can see math is not my strongest subject

OpenStudy (anonymous):

Remember, that number you found is b^2. You need to take the square root of 14.75 in order to find b.

OpenStudy (anonymous):

so 3.841^2

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Thank you so much for your help. It is greatly appreciated

OpenStudy (anonymous):

You're welcome.